If a probability that A,B and C will solve a problem is 1/2,3/4 and2/3.what is the probability that the problem will be solved



1
Expert's answer
2021-03-05T00:52:00-0500

Find the probability Q that the problem will not be solved:

"Q = (1 - p(A))(1 - p(B))(1 - p(C)) = \left( {1 - \frac{1}{2}} \right)\left( {1 - \frac{3}{4}} \right)\left( {1 - \frac{2}{3}} \right) = \frac{1}{{24}}"

Then the wanted probability, as the probability of the opposite event, is

"P = 1 - Q = 1 - \frac{1}{{24}} = \frac{{23}}{{24}}"

Answer: 23/24